OpenStudy (anonymous):

\[\int\frac{1}{1+x^{4}}dx\]

OpenStudy (anonymous):

Without integration by partial fractions.

OpenStudy (lgbasallote):

can i give it now? :S

OpenStudy (anonymous):

Go ahead. :P

OpenStudy (lgbasallote):

ill wait some more :p

OpenStudy (anonymous):

naah, noow; just see what you got :P

OpenStudy (anonymous):

is it trigonometric substitution?

OpenStudy (anonymous):

You can manipulate it; not sure if trig-sub would work. I am not sure if \(x^2=tan\theta\) would work. @lgbasallote: Can tell us his method. :P

OpenStudy (lgbasallote):

ill let @Mimi_x3 express his/her methods first

OpenStudy (anonymous):

lol, who is Mimi_x3?

OpenStudy (apoorvk):

No one remembered me I see. :\ Still. Divide both numerator and denominator by x^2

OpenStudy (anonymous):

trig sub get messy...:(

OpenStudy (anonymous):

yeah, it does; but the manipulation gets messier i suppose :/

OpenStudy (apoorvk):

\[\int \frac{\frac1{x^2}}{\frac1{x^2} + x^2}.dx\] \[=\int \frac{\frac1{x^2}}{(\frac1{x} +x)^2-2}.dx\] Now, let "x + 1/x" be t the, (-1/(x^2)).dx = dt So, \[\int \frac{-dt}{t^2 -2}\] INTEGRATE THIS!!!!

OpenStudy (apoorvk):

The loving integrand formula for this one is:\[\int\frac{dx}{x^2 - a^2}= \frac{1}{2a}ln(\frac{x-a}{x+a}) + c\]

OpenStudy (anonymous):

Nice manipulation; i havent though of that. \[\frac{1}{2} \int\limits\frac{2}{1+x^{2}} dx => \frac{1}{2} \int\limits\frac{(x^{2}+1)-(x^{2}-1)}{1+x^{4}} \] this is another way (:

OpenStudy (apoorvk):

Yeah, this is another way, but then you have two integrals, so a wee bit messier, but nonetheless , a good one :)

OpenStudy (anonymous):

I made a typo. xD \[\frac{1}{2} \int\limits\frac{2}{1+x^{4}} dx\]

OpenStudy (experimentx):

another way \[ 1 + x^4 = (1 + ix^2)(1 - ix^2)\]

OpenStudy (anonymous):

lol, complex numbers..

OpenStudy (experimentx):

if you are comfortable with complex numbers ...

OpenStudy (anonymous):

yeah, abit harder i suppose..

OpenStudy (callisto):

Hold on.... Let t = x + 1/x dt = 1 - 1/x^2 dx .... Isn't it? (Sorry, I haven't done differentiation and integration for long..)

OpenStudy (callisto):

I mean ... if dt = 1 - 1/x^2 dx is true, then (-1/(x^2)).dx = dt is not correct... Then, we cannot get \(\int \frac{-dt}{t^2 -2}\)

OpenStudy (apoorvk):

@Callisto you are right, I blooped big time. :/ Thanks... Actually i should have done it this way: \[=\int \frac{\frac1{x^2}}{(x -\frac1{x} )^2+2}.dx\] Now it should work out like this: Let x - 1/x be 't' Then (1/(x^2)).dx = dt So, \[\int \frac{dt}{t^2 + 2}dt\] And this one's even easier. \[\int \frac{dx}{x^2+a^2} = \frac 1 a \tan^{-1}\frac x a + c\]

OpenStudy (callisto):

No... I meant Let x - 1/x be 't' Then (1/(x^2)).dx = dt <- but isn't it dt = 1+ 1/x^2 dx ?

OpenStudy (apoorvk):

I AM SO SO SOOOOOOOOOOOO STUPID!!!! Take back those virtual accolades, i don't deserve them! :(

OpenStudy (anonymous):

lol, forget about this integral :)

OpenStudy (callisto):

No... The method is good. But the question is still unsolved..

OpenStudy (anonymous):

Don't worry about it (: I already gave you a method that would work.

OpenStudy (apoorvk):

Okay, i seem to have got his now. \[=\int \frac{\frac1{x^2}}{x^2 +\frac1{x^2} }.dx =\frac 1 2 ( \int \frac{1+\frac1{x^2}}{(x -\frac1{x} )^2+2}.dx - \int \frac{1 - \frac1{ x^2} }{(x +\frac1{x} )^2-2}.dx) \] Hmm should work now.

OpenStudy (apoorvk):

*this lool.

OpenStudy (apoorvk):

I hope I didn't do anything stupid again. -_-

OpenStudy (apoorvk):

Okay let's work this thing out, just to be sure: In the first part, let \[x - \frac 1 x = t\]\[or, (1 + \frac 1 {x^2})dx = dt\] In the second part, \[Let~~ x +\frac 1 x = m\]\[So,~~ (1 - \frac 1 {x^2})dx = dm\] So, what we have is: \[\frac 1 2 (\int \frac{dt}{t^2 + 2} - \int \frac{dt}{m^2 - 2})\] \[=\frac 1 2 ( \frac 1 {\sqrt 2}tan^{-1}(\frac t {\sqrt 2}) - \frac 1 {2\sqrt 2}ln(\frac{m-\sqrt2}{m +\sqrt2}) + k)\] Now substitute t = x - 1/x and m = x + 1/x, and enjoy!!

OpenStudy (anonymous):

@Callisto: Another method. \[\frac{1}{2} \int\limits\frac{2}{1+x^{4}} dx => \frac{1}{2} \int\limits\frac{(x^{2}+1)-(x^{2}-1)}{1+x^{4}} =>\] \[\frac{1}{2} \int\limits\frac{x^{2}+1}{1+x^{4}} -\frac{x^{2}-1}{1+x^{4}} \] I think that you're able to do it now.